OpenStudy (anonymous):

find the exact area of the surface obtained by rotating the given curve about the x-axis.

OpenStudy (anonymous):

x = cos^3 t y= sin^3 t 0<theta<pi/2

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

@Loser66 can you help me with this problem

OpenStudy (anonymous):

i am confused as to what to do

OpenStudy (loser66):

@freckles

OpenStudy (freckles):

we want to use this \[\int\limits_{}^{} 2 \pi y \sqrt{(x')^2+(y')^2} dt \]

OpenStudy (anonymous):

alright

OpenStudy (freckles):

\[\int\limits_{0}^{\frac{\pi}{2}}2\pi \sin^3(t) \sqrt{((\cos^3(t))')^2+((\sin^3(t))')^2} dt\]

OpenStudy (freckles):

we need to differentiate the cos cos^3(t) and the sin^3(t) thingy w.r.t to t

OpenStudy (freckles):

for example \[\frac{d}{dt} \sin^3(t)=3\sin^2(t) \cdot \cos(t)\] then square both sides to find the square of the derivative since our formula calls for that: \[(\frac{d}{dt} \sin^3(t))^2=(3 \sin^2(t) \cos(t))^2 =9 \sin^4(t) \cos^2(t)\] so you try both of these things for the cos^3(t) part

OpenStudy (anonymous):

 this is what i have so far

OpenStudy (freckles):

\[\frac{d}{dt}\cos^3(t)=3\cos^2(t)(-\sin(t)) \\ (\frac{d}{dt} \cos^3(t))^2=9\cos^4(t)\sin^2(t)\] so let's see we are going to put this in and it looks like you have it ... \[\int\limits_{0}^{\frac{\pi}{2}} 2\pi \sin^3(t) \sqrt{9\cos^4(t)\sin^2(t)+9\sin^4(t)\cos^2(t)} dt\] well except that -1 part

OpenStudy (freckles):

(-1)^2=1

OpenStudy (freckles):

now we could factor inside that square root part

OpenStudy (anonymous):

how do we do that